How much money do people spend on graduation gifts? In 2007, a federation surveyed 2815 consumers who reported that they bought one or more graduation gifts that year. The sample was selected in a way designed to produce a sample representative of adult Americans who purchased graduation gifts in 2007. For this sample, the mean amount spent per gift was $52.65. Suppose that the sample standard deviation was $20. Construct a 98% confidence interval for the mean amount of money spent per graduation gift in 2007. (Round your answers to three decimal places.)
Answer #1
sample mean 'x̄= 52.650
sample size   n= 2815.00
sample std deviation s= 20.000
std error 'sx=s/√n= 0.3770
for 98% CI; and 2814 df, value of t= 2.328
margin of error E=t*std error    = 0.877
lower bound=sample mean-E = 51.773
Upper bound=sample mean+E = 53.527
from above 98% confidence interval for population mean =(51.773 ,53.527)